# signature_mor.bib

### v1

@@ 1 Introduction
The path signature (Chen [chen1957integration]) – i.e., the sequence of all iterated integrals of the components of a path $x:[0,T]\to\mathbb{R}^{d}$ – plays a fundamental rule in the analysis of (rough) paths and functions on path space. On the theoretical side, the signature is the fundamental building block of Lyons’ theory of rough paths [lyons1998differential], providing pathwise analysis of differential equations driven by rough paths (i.e., Hölder continuous paths of arbitrary Hölder coefficient), see [friz2020course, friz2010multidimensional] for modern accounts of the theory. Indeed, a rough path is nothing but a path together with a (properly truncated) signature.
@@
In applications, signatures play a fundamental role for approximations of functionals on (rough) path space, comparable to polynomials on finite-dimensional spaces. Indeed, it is well-known that linear functionals on signatures are universal approximators, i.e., that continuous functions mapping path-space to the real numbers can be approximated by linear functionals of the signature when restricted to compact sets. Several alternative formulations also exist, we refer to [lee2023signature] for a recent review article and [bayer2023primal] for a version on approximation in an $L^{p}$-sense.
@@
Of course, similar approximation results are also available in a probabilistic rather than pathwise setting, i.e., when considering stochastic processes rather than single paths. A corresponding universal approximation result in terms of signatures (of semi-martingales) can be found in [cuchiero2022universal], but we also like to mention the classical stochastic Taylor expansion in terms of iterated integrals of the Brownian motion.
@@
The path signature has been found to be a powerful feature map in machine learning, when dealing with time series data, see T. Lyons’ ICM lecture [lyons2014rough], for instance, [perez2018signature, buehler2020generating, morrill2020generalised, lyons2022signature, arrubarrena2024novelty] for specific examples. Here, both the fact that signatures provide (efficient) encodings of paths (of arbitrary length) in a fixed vector space and the invariance under time re-parameterizations are crucial.
@@
Beyond machine learning, signatures can be used as a fundamental building block for modelling and stochastic analysis of processes with memory – or, more generally, when we only want to impose minimal assumptions on the model. An example for the use of signatures in modelling is [cuchiero2022signature], where a diffusion-type model with coefficients given as linear functionals of the signature of the underlying Brownian motion is fitted to market prices of options.
@@
Beyond modelling, the signature can be a powerful tool for analysis and numerics of non-Markovian processes. In particular, for BSDEs (see, for instance, [BFZ24]) or stochastic optimal control ([KLPA20, bayer_hager_riedel_book24]) we are often dealing with computations of conditional expectations w.r.t. the relevant sigma-algebra $\mathcal{F}_{t}$. In the Markovian case, this corresponds to a finite-dimensional regression problem. But in the non-Markovian case (i.e., with memory), such conditionals expectations generally depend on the whole history of the driving path up to time $t$. Such path-dependent conditional expectations can be efficiently computed as (linear or non-linear) functions of the signature of the underlying path, see [bayer2023primal].
@@
From a computational point of view, the signature poses severe challenges mainly due to its size. Indeed, a priori, the signature – the collection of all iterated integrals – is an infinite-dimensional object, and, hence, not directly computable. For some applications, the signature kernel (see [lee2023signature]) can be used instead, which is actually computable as a solution of a PDE. Otherwise, the most natural approach is to truncate the signature, i.e., to restrict the computations to iterated integrals up to a given degree. Even so, the size of the (finite) sequence of iterated integrals of degree up to $m$ increases rapidly in terms of $m$ as well as the dimension $d$ of the underlying space: in fact, it is easy to see that there are $1+d+\cdots+d^{m}$ such terms – including one trivial term of order 00, think of an “intercept” in a linear model.
@@
This work addresses these concerns by exploring model order reduction (MOR) techniques specifically tailored for signature-based models. MOR has a long tradition in the field of deterministic control systems [moo1981], where difficult to control and hard to observe states are neglected to receive a low-dimensional approximation of a large-scale model. This work is the foundation for establishing balancing-related MOR for controlled SDEs [beckerhartmann, bennerdamm, redmannbenner, bennerdammcruz]. Complexity reduction in such probabilistic frameworks has an enormous impact, e.g., when sampling methods are applied. However, this work relies on certain stability assumptions that are not necessarily satisfied in practice. A way to overcome this issue is to use the ansatz of [morGawJ90] established for controlled systems of ordinary differential equations and extended to SDEs in [redmann_jamshidi]. The control theory perspective on MOR does no longer work when dimension reduction is applied to financial models. Let us refer to [irka_finance, mor_heston], where a bridge was built between the previously mentioned concepts and complexity reduction for large-scale asset price models. Detached from control concepts and stability assumptions, MOR for signatures is studied in this paper. We develop a robust mathematical framework to reduce the complexity of these models while preserving their essential features, provide a comprehensive method for deriving reduced order models of truncated signatures, with discussions on their theoretical foundations and practical implications. The efficiency of the proposed techniques is illustrated through numerical examples, including an application to the (rough) Bergomi model–a widely used model in financial markets. The results demonstrate the effectiveness of MOR in reducing the dimension without a significant loss of accuracy.
@@
Indeed, we apply MOR to realistic signature-models of financial markets, as derived from popular equity models (the classical Bergomi model [bergomi2015stochastic] and the rough Bergomi model [bayer2016pricing], respectively). Starting from signature models of dimension $n=1365$ in the Bergomi case, we find that a reduced model of dimension $\widetilde{n}=5$ achieves a relative error of order $10^{-2}$ in the implied volatility of option prices, $\widetilde{n}=11$ is sufficient for a relative error of $10^{-4}$, and $\widetilde{n}=27$ is exact up to machine precision, see Figure 4. We observe that the same order of accuracy is obtain in the $L^{2}$-sense, see Figure 2. This means that our reduced system yields a strong approximation.
@@
We start by providing an introduction to path signatures in Section 2. The basic structures are introduced in the setting of smooth paths and later extended to signatures of continuous semimartingales. We also provide the fundamental approximation results for signatures and minimal necessary results from rough path theory. As a motivation, we summarize results on signature models in finance by [cuchiero2022signature].
Section 3 provides a derivation of MOR for signatures. Note that (truncated) signatures of continuous semi-martingales solve a linear system of SDEs with specific structure, to which general MOR for SDEs is adapted. The reduced signature equation is introduced in Section 3.3.
@@ 2 Signatures
2.1 Signatures of smooth paths
(2.1) $\mathbbm{x}^{I}_{s,t}\coloneqq\int_{s<t_{1}<\cdots<t_{n}<t}\dot{x}^{i_{1}}_{t_{1}}\mathrm{d}t_{1}\cdots\dot{x}^{i_{n}}_{t_{n}}\mathrm{d}t_{n}\eqqcolon\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x^{i_{1}}_{t_{1}}\cdots\mathrm{d}x^{i_{n}}_{t_{n}}\in\mathbb{R},$
where $0\leq s\leq t\leq T$. Here, $x^{i}_{t}$ denotes the $i$th coordinate of the path $x$ evaluated at time $t$ for a fixed basis $e_{1},\ldots,e_{d}$ of $V$. We prefer to use the notation “$\mathrm{d}x_{t}$” rather than “$\dot{x}_{t}\mathrm{d}t$” because the former corresponds to the standard notation in the non-smooth case to be considered later – think of paths of Brownian motion.
(2.2) $\mathbbm{x}^{<\infty}_{s,t}\coloneqq 1+\sum_{n=1}^{\infty}\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x_{t_{1}}\otimes\cdots\otimes\mathrm{d}x_{t_{n}}\in T((V)),$
with the extended tensor algebra $T((V))$ being defined as
(2.3) $T((V))\coloneqq\prod_{n=0}^{\infty}V^{\otimes n}.$
Here, the initial term “$1$” is considered as the (sole) entry in $V^{\otimes 0}\simeq\mathbb{R}$. Note that elements of the extended tensor algebra $T((V))$ have infinitely many terms, and are, hence, comparable to formal power series – in $d$ non-commuting variables $e_{1},\ldots,e_{d}$. For numerical purposes, signatures need to be truncated at a finite degree. Hence, we also define
(2.4) $\mathbbm{x}^{\leq m}_{s,t}\coloneqq 1+\sum_{n=1}^{m}\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x_{t_{1}}\otimes\cdots\otimes\mathrm{d}x_{t_{n}}\in T^{m}(V)\coloneqq\bigoplus_{n=0}^{m}V^{\otimes n}.$
Both $T((V))$ and $T^{m}(V)$ are algebras under the tensor product $\otimes$ – the product of formal power series in non-commuting variables. We refer to [bayer2021optimal] for more information. Using this product, we see that the signature (formally) solves a controlled ODE, namely $\dot{\mathbbm{x}}_{0,t}^{<\infty}=\mathbbm{x}_{0,t}^{<\infty}\otimes\dot{x}_{t}$ or (in a notation easier to adapt for non-smooth paths)

### v2

# Dimension reduction for path signatures

@@ 1 Introduction
The path signature (Chen [16]) – i.e., the sequence of all iterated integrals of the components of a path $x:[0,T]\to\mathbb{R}^{d}$ – plays a fundamental role in the analysis of (rough) paths and functions on path space. On the theoretical side, the signature is the fundamental building block of Lyons’ theory of rough paths [31], providing pathwise analysis of differential equations driven by rough paths (i.e., Hölder continuous paths of arbitrary Hölder coefficient), see [22, 23] for modern accounts of the theory. Indeed, a rough path is nothing but a path together with a (properly truncated) signature.
@@
In applications, signatures play a fundamental role for approximations of functionals on (rough) path space, comparable to polynomials on finite-dimensional spaces. Indeed, it is well-known that linear functionals on signatures are universal approximators, i.e., that continuous functions mapping path-space to the real numbers can be approximated by linear functionals of the signature when restricted to compact sets. Several alternative formulations also exist, we refer to [29] for a recent review article and [6] for a version on approximation in an $L^{p}$-sense.
@@
Of course, similar approximation results are also available in a probabilistic rather than pathwise setting, i.e., when considering stochastic processes rather than single paths. A corresponding universal approximation result in terms of signatures (of semimartingales) can be found in [19], but we also like to mention the classical stochastic Taylor expansion in terms of iterated integrals of the Brownian motion.
@@
The path signature has been found to be a powerful feature map in machine learning, when dealing with time series data, see T. Lyons’ ICM lecture [32], for instance, [35, 15, 34, 30, 1] for specific examples. Here, both the fact that signatures provide (efficient) encodings of paths (of arbitrary length) in a fixed vector space and the invariance under time re-parameterizations are crucial.
@@
Beyond machine learning, signatures can be used as a fundamental building block for modelling and stochastic analysis of processes with memory – or, more generally, when we only want to impose minimal assumptions on the model. An example for the use of signatures in modelling is [18], where a diffusion-type model with coefficients given as linear functionals of the signature of the underlying Brownian motion is fitted to market prices of options.
@@
Beyond modelling, the signature can be a powerful tool for analysis and numerics of non-Markovian processes. In particular, for BSDEs (see, for instance, [8]) or stochastic optimal control ([28, 5]) we are often dealing with computations of conditional expectations w.r.t. the relevant sigma-algebra $\mathcal{F}_{t}$. In the Markovian case, this corresponds to a finite-dimensional regression problem. But in the non-Markovian case (i.e., with memory), such conditionals expectations generally depend on the whole history of the driving path up to time $t$. Such path-dependent conditional expectations can be efficiently computed as (linear or non-linear) functions of the signature of the underlying path, see [6].
@@
From a computational point of view, the signature poses severe challenges mainly due to its size. Indeed, a priori, the signature – the collection of all iterated integrals – is an infinite-dimensional object, and, hence, not directly computable. For some applications, the signature kernel (see [29]) can be used instead, which is actually computable as a solution of a PDE. Otherwise, the most natural approach is to truncate the signature, i.e., to restrict the computations to iterated integrals up to a given degree. Even so, the size of the (finite) sequence of iterated integrals of degree up to $m$ increases rapidly in terms of $m$ as well as the dimension $d$ of the underlying space: in fact, it is easy to see that there are $1+d+\cdots+d^{m}$ such terms – including one trivial term of order 0, think of an “intercept” in a linear model.
This work addresses these concerns by exploring model order reduction (MOR) techniques specifically tailored for signature-based models. MOR has a long tradition in the field of deterministic control systems [33], where difficult to control and hard to observe states are neglected to receive a low-dimensional approximation of a large-scale model. This work is the foundation for establishing balancing-related MOR for controlled SDEs [9, 12, 13, 11]. Complexity reduction in such probabilistic frameworks has an enormous impact, e.g., when sampling methods are applied. More specific to financial models, an important potential use case of MOR is derivative pricing, requiring many simulations based on one (calibrated) asset price model. For details on the application of MOR to derivative pricing, including in the case of American options, see [36].
@@
However, the above mentioned works rely on certain stability assumptions that are not necessarily satisfied in practice. A way to overcome this issue is to use the ansatz of [24] established for controlled systems of ordinary differential equations and extended to SDEs in [37]. The control theory perspective on MOR does no longer work when dimension reduction is applied to financial models. Let us refer to [38, 36], where a bridge was built between the previously mentioned concepts and complexity reduction for large-scale asset price models. Detached from control concepts and stability assumptions, MOR for signatures is studied in this paper. We develop a robust mathematical framework to reduce the complexity of these models while preserving their essential features, provide a comprehensive method for deriving reduced order models of truncated signatures, with discussions on their theoretical foundations and practical implications. The efficiency of the proposed techniques is illustrated through numerical examples, including an application to the (rough) Bergomi model–a widely used model in financial markets. The results demonstrate the effectiveness of MOR in reducing the dimension without a significant loss of accuracy.
@@
Indeed, we apply MOR to realistic signature-models of financial markets, as derived from popular equity models (the classical Bergomi model [14] and the rough Bergomi model [3], respectively). Starting from signature models of dimension $n=1365$ in the Bergomi case, we find that a reduced model of dimension $\widetilde{n}=5$ achieves a relative error of order $10^{-2}$ in the implied volatility of option prices, $\widetilde{n}=11$ is sufficient for a relative error of $10^{-4}$, and $\widetilde{n}=27$ is exact up to machine precision, see Figure 4. We observe that the same order of accuracy is obtain in the $L^{2}$-sense, see Figure 2. This means that our reduced system yields a strong approximation.
@@
We start by providing an introduction to path signatures in Section 2. The basic structures are introduced in the setting of smooth paths and later extended to signatures of continuous semimartingales. We also provide the fundamental approximation results for signatures and minimal necessary results from rough path theory. As a motivation, we summarize results on signature models in finance by [18].
Section 3 provides a derivation of MOR for signatures. Note that (truncated) signatures of continuous semimartingales solve a linear system of SDEs with specific structure, to which general MOR for SDEs is adapted. The reduced signature equation is introduced in Section 3.3.
@@ 2 Signatures
Signatures can be seen as part of the repertoire of mathematical finance by now. For this reason, we only give a very cursory introduction here, mainly to fix notations. For readers not yet familiar with signatures, a detailed exposition is given in Appendix A.
Given a finite-dimensional real vector space $V$ and $T>0$, let $x:[0,T]\to V$ be an $\alpha$-Hölder continuous path ($\alpha\in(0,1]$), and denote by $\mathbf{x}\in\mathscr{C}^{\alpha}_{g}([0,T];V)$ a geometric rough path lift of $x$. We will often assume that the first component $x^{1}_{t}\equiv t$, and denote the set of geometric $\alpha$-Hölder rough paths satisfying this condition by $\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)$.
The (truncated) signature of $\mathbf{x}$ is the collection of all iterated integrals of the form
$\int_{s<t_{1}<\cdots<t_{n}<t}\mathrm{d}x^{i_{1}}_{t_{1}}\cdots\mathrm{d}x^{i_{n}}_{t_{n}},\ i_{1},\ldots,i_{n}\in\{1,\ldots,\dim V\},\ n\in\mathbb{N},\ 0\leq s\leq t\leq T,$
(up to order $m\in\mathbb{N}$) and is denoted by $\mathbbm{x}^{<\infty}_{s,t}$ ($\mathbbm{x}^{\leq m}_{s,t}$, respectively) for $0\leq s\leq t\leq T$. It takes values in the extended tensor algebra $T((V))\coloneqq\prod_{n=0}^{\infty}V^{\otimes n}$ or its truncated version $T^{m}(V)\coloneqq\prod_{n=0}^{m}V^{\otimes n}$.
In many respects, signatures, or, more precisely, linear functionals on the signature tensor, can be seen as natural generalizations of polynomials from Euclidean space to path-space. In particular, the signature $\mathbbm{x}^{<\infty}_{0,T}$ (essentially) characterizes the underlying path $x$, and the space of linear functionals on the signature form an algebra of functions on path space. As a consequence, the linear functionals on the signature are dense in the continuous functions on path-space on compacts – see Theorem A.9 for a precise statement. Hence, signatures provide universal features on path space, and can be seen as natural building blocks for approximation of continuous functions on path-space, similar to the role of polynomials in Euclidean space.
A particularly important property of signatures from the point of view of MOR is that it is given as the solution of a linear system of (controlled) differential equations. Indeed, we formally have